Joe bought 8 sacks of birdseed for his petshop. Each sack weighed 15kg. He repacked the seed into 450g bags.
1. How many full bags did he pack?
2. Joe bought the birdseed at 20$ per sack. How much profit does he make if he sells the bags for 1$ each?

Answer by edjones(8007)

1)
8*15000=120000
120000/450=266 full bags.
.
2)
20*8=$160 Cost
266-160=$106 Profit
.
Ed